OpenStudy (anonymous):

Tommy can paint a fence in 10 hours. If his friend huck helps, they can paint the fence in one hour. How fast would it take huck to paint the fence by himself?

OpenStudy (saifoo.khan):

I think there is a mistake here.

OpenStudy (saifoo.khan):

1/10 + 1/x = 1 Solve for x.

OpenStudy (anonymous):

hmm... this is depressing, I still cant figure it out...

OpenStudy (saifoo.khan):

\[\frac{1}{10}+\frac 1x = \frac 11\]

OpenStudy (anonymous):

Since Tommy alone can paint the full fence in 10 hours It means that in 1 hour Tommy can paint 1/10 of the fence Since Tommy and Huck together finish the work together in 1 hour => Huck can finish 9/10 of the fence in 1 hour (because 1/10 is done by Tommy) So if Huck can finish 9/10 of fence in one hour , then he can finish the full fence in time 1 1 10 10 = ------- = ---- * ----- = ----- hours 9 1 9 9 ----- 10